What is the slope of the v-t graph if average velocity is equal to instantaneous velocity? Is it 0 or can it be something else also?
Can a body have acceleration with zero velocity?please give an example.
Thanks
Jai

Grade:12

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,

ds/dt=s/t
=>lns=lnt+lnc
=>s=c1*e^t

also
yes
particle projected at maxmium height has zero verticle velocity but zero verticle acceleration.